OpenStudy (anonymous):

***CHECK*** Simplify and Identify the domain. -4/4x+8*5(x+3)/x^2-9 A. -5/(x+2)(x-3),(x does not = -2,3) B. -5/(x+2)(x-3), (x does not = -3, -2,3) C. -5/(x+2)(x-3), (x does not = 3) D. -1/8x^2, (x does not = -3, -2,3)

OpenStudy (anonymous):

I think A

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Can you write the expression again? I don't know what the fractions are.

OpenStudy (anonymous):

-4/4x+8*5(x+3)/x^2-9 {-4/4x+8*5(x+3} over {x^2-9}

OpenStudy (igreen):

\(\sf \dfrac{\dfrac{-4}{4x} + 8 \times 5(x + 3)}{x^2 - 9}\)?

OpenStudy (igreen):

Is that it?

OpenStudy (igreen):

Is it (-4)/(4)x or is it (-4)/(4x)

OpenStudy (anonymous):

The slash is wrong my bad /.\ -4/4x+8*5(x+3)*x^2-9

OpenStudy (igreen):

Oh, well, you still need to explain the first division sign

OpenStudy (anonymous):

\[\frac{ -4 }{ 4x+8 }\times \frac{ 5(x+3) }{ x^2-9 }\]

OpenStudy (igreen):

A is incorrect..

OpenStudy (igreen):

You're correct about x not being -2 and 3, but there's one more.

OpenStudy (anonymous):

Ahh so B

OpenStudy (anonymous):

@iGreen ?

OpenStudy (igreen):

Yep!

OpenStudy (anonymous):

Thx! c: